# Superconductivity in spin-orbit coupled BaBi3 formed by in situ reduction of bismuthate films

### v1

@@ Abstract
Oxygen-scavenging at oxide heterointerfaces has emerged as a powerful route for stabilizing metastable phases that exhibit interesting phenomena, including high-mobility two-dimensional electron gases and high Tc superconductivity. We investigate structural and chemical interactions at the heterointerface formed between Eu metal and the charged-ordered insulator, BaBiO3, which leads to emergent superconductivity at 6 K. A combination of X-ray diffraction and electron microscopy measurements shows that oxygen scavenging by the Eu adlayer leads to the formation of EuOx and the superconducting intermetallic BaBi3. The superconducting state is quasi-two-dimensional, as evidenced by the angle-dependent magnetotransport measurements and current-voltage characteristics. An in-plane upper critical field Bc2 significantly exceeding the Pauli paramagnetic limit is observed, possibly reflecting features of Ising superconductivity. The strong spin-orbit coupling at the Bi sites may pave the way for the realization of high-Tc topological superconductivity.
@@ I Introduction
Interest in Bi-rich superconductors continues due to the potential role of the strong spin-orbit coupling of Bi-6p electrons in enhancing superconductivity, and magnetic field-induced transitions from conventional s-wave to unconventional nodal d-wave superconductivity [9, 10]. Intermetallic ABi3 compounds (where A = Ba or Sr) have received attention because they exhibit superconductivity enhanced by spin-orbit coupling[11, 12, 9, 13]. Theoretical studies show that spin-orbit coupling suppresses strong Fermi nesting and increases electron-phonon coupling matrix elements associated with superconductivity[9]. At ambient pressure, BaBi3 crystallizes in a tetragonal structure (P4/mmm, a=5.06 , c=5.13 )[13, 12, 14, 15], featuring corner-sharing Bi6 octahedra (Figure 1B). Magnetotransport measurements on single crystals of BaBi3 and SrBi3 reveal superconducting transitions with onset temperatures of approximately 5.9 K and 5.6 K, respectively [12, 13]. Furthermore, scanning tunneling microscopy (STM) studies demonstrate the effect of symmetry breaking and disorder on the superconducting state in BaBi3 epitaxial films, grown in situ by molecular beam epitaxy[16]. However, to date, there have been no reports on the magnetotransport properties of thin BaBi3, due to its high air sensitivity [13].
Here, we report the emergence of superconductivity at approximately 6 K in nominally Eu-capped p-BBO thin films grown via molecular beam epitaxy. X-ray diffraction measurements reveal that oxygen scavenging by the in-vacuo deposited metallic Eu overlayer leads to the reduction of the monoclinic p-BBO layers into tetragonal superconducting BaBi3 and rock-salt BaO. Several measurements, including angle-dependent magnetotransport and current-voltage characteristics, confirm the quasi-two-dimensional nature of the superconductivity. An in-plane upper critical field Bc2 far exceeding the Pauli paramagnetic limit, raising intriguing possibilities of spin-protected superconducting states, such as Ising superconductivity[17]. These findings highlight the potential of utilizing in situ reduction of bismuthate heterostructures as a platform for exploring novel quantum states. The strong spin-orbit coupling at Bi sites, may pave the way for the realization of high-temperature topological superconductivity, a long-sought phenomenon in perovskite-based systems[18, 19].
@@
Synthesis:
The samples were fabricated in a custom-oxide molecular beam epitaxy system. Bi, Ba, Eu, and Al were evaporated from effusion cells, with their growth rates calibrated by a quartz crystal monitor (QCM). p-BBO thin films of thicknesses ranging from 12 nm (30 unit cells) to 36 nm (90 unit cells) were grown on (001)-oriented SrTiO3 (STO) substrates at a substrate temperature of 585∘ C in an oxygen plasma with an oxygen partial pressure of 6.5x10-6 Torr and a plasma power of 250 W. Sharp reflection high-energy electron diffraction (RHEED) diffraction streaks observed during p-BBO growth shown in Figure 2A, are indicative of atomically flat p-BBO layers. The in-plane lattice parameter determined from the RHEED spacings is 4.35 , in agreement with bulk stoichiometric p-BBO [5]. After growth, the p-BBO films were cooled to 500∘ C for the deposition of Eu metal in vacuo. RHEED patterns measured during Eu growth (Figure 2A) are consistent with the formation of cubic EuO from the topotactic reduction of the p-BBO layer [20, 21, 22]. The films were capped with 3 nm AlOx at room temperature to prevent oxidation when exposed to ambient conditions.
@@
Figure 2B shows a comparison of specular scans along the (00L) direction for the nominal Eu/p-BBO sample and a single-component p-BBO film on STO. For the single-component p-BBO film, Bragg peaks are observed for the STO substrate and stoichiometric p-BBO. The calculated out-of-plane lattice constant for the p-BBO film is determined to be 4.36 in agreement with the measured values for the monoclinic bulk phase [5]. In contrast to the single-component p-BBO film, no p-BBO Bragg peaks were observed for the nominal Eu/p-BBO sample. Bragg peaks are observed at 32.316o and 34.997o corresponding to the (002) Bragg peaks of rock-salt BaO and face-centered tetragonal BaBi3, respectively. The calculated out-of-plane lattice constants are 5.52 and 5.11 in good agreement with the bulk lattice constants for BaO (5.53 ) and BaBi3 (5.13 ) [23, 12]. The lattice constant of NaCl-type EuO is 5.1439 ,[24] hence, the EuO (002) Bragg peak will overlap with the BaBi3 (002) peak. To confirm the formation of BaBi3, we deposit Al metal directly on a p-BBO film at 500∘C. We find that the Al metal effectively reduces the p-BBO film to form BaBi3 and BaO as evidenced by XRD measurements in Figure 2B.
@@
Scanning transmission electron microscopy (STEM) measurements were performed on the films to further investigate the structural properties of the heterostructures. Figure 2C shows a high-angle annular dark field (HAADF) STEM image of a cross-section of a nominal Eu/p-BBO film on STO. Sharp interfaces are observed between the nominal layers, indicative of insignificant intermixing. The energy-dispersive X-ray spectroscopy (EDX) profiles of the section are shown in Figure 2D. The EDX maps reveal significant chemical segregation in the nominal p-BBO region. The nominal Eu and STO regions are oxygen-rich, while the nominal p-BBO section is oxygen-poor, consistent with the reduction of p-BBO by the Eu adlayer. Ba-rich regions are observed at the nominal p-BBO/STO and Eu/p-BBO interfaces. In the nominal p-BBO region, we observe lateral phase-separated Ba and Bi-rich regions, as shown in Figure S1. The oxygen content in the Ba-rich regions is significantly higher than in the Bi-rich sections, consistent with the BaO and BaBi3 Bragg peaks observed in the XRD measurements.
Transport Measurements:
Next, we investigate the transport properties of nominal Eu/p-BBO heterostructures from 1.8 - 300 K. Figure 3A shows the temperature-dependent resistance of a nominal Eu (16 nm)/p-BBO (36 nm) heterostructure (Sample 1). As the temperature decreases to 7 K, the resistance decreases, exhibiting characteristic metallic behavior. Upon further cooling, the sample undergoes a transition to a superconducting state with a narrow transition width of approximately 0.133 K, reaching a zero resistance state at 5.77 K (). The onset superconducting temperature (defined as the temperature at which the resistance drops to 90% of its normal state value), is 6.03 K, significantly higher than the previously reported value of 3.5 K for EuO/BaBiO3 heterostructures[20] but in agreement with the reported values for BaBi3[12, 13, 9, 14, 15]. Above Tc, the resistance exhibits a linear temperature dependence, a feature not discussed in bulk BaBi3 samples. This linearity is more clearly demonstrated by the linear fitting of the resistance vs. temperature data shown in Figure S2. Such behavior is indicative of strange-metal characteristics commonly observed in unconventional superconductors [25]. The inset of Figure 3A shows the temperature-dependent resistance of sample 1 with a magnetic field (), applied out of the plane. is suppressed as increases, further confirming superconductivity. Magnetoresistance curves with at different temperatures close to the superconducting transition are plotted in Figure 3B. In the normal state above Tc, the resistance exhibits a quadratic dependence on the applied magnetic field, similar to conventional metals[26]. In particular, below Tc, a resistance peak appears above the critical magnetic field, suggesting disorder arising from granularity[27].
@@
The temperature-dependent Hall resistance exhibits a linear dependence on the magnetic field, B, as shown in Inset Figure 3C. This linear behavior confirms the presence of a single type of charge carrier. However, a change in slope and, consequently, the sign of the Hall coefficient occurs at 65 K, which can be seen more clearly in Figure 3C. This indicates a transition from low-temperature electronic conduction to hole-dominant conduction at high temperatures. The sheet carrier density and mobility at 10 K are 1.6 1016 cm-2 and 1.2 cm2/Vs, respectively. These carrier density values are consistent with previous observations[20]. The carrier concentrations are 3-4 orders of magnitude higher than observed for the two-dimensional electron gas at SrTiO3 [28, 29] and KTaO3 [30, 31, 32, 33] interfaces, suggesting that the carriers may extend through the film thickness (). Furthermore, Figure 3D compares the resistance as a function of the nominal p-BBO thickness. The number of Eu monolayers is tuned to match the p-BBO thicknesses. The transition temperatures are comparable and independent of the nominal p-BBO thicknesses.
@@
The critical fields for both out-of-plane, , and in-plane configurations were determined using the 90% normal resistance criterion. and are plotted as a function of temperature in Figure 4D. To determine the upper critical fields at T = 0 K, the data were fitted using the two-dimensional Ginzburg-Landau (GL) model for the out-of-plane configuration [34]:
@@
where represent the extrapolated critical fields at 0 K. The extracted value of is 6.24 T, exceeding the reported value of 2.2 T for BaBi3 single crystals [12, 35]. The temperature variation of the in-plane upper critical field deviates from the 2D Ginzburg-Landau (GL) model for the parallel configuration, exhibiting a higher value at T= 0 K as suggested by the dashed line in Figure 4D. The disparity between the in-plane and out-of-plane upper critical fields confirms the anisotropic nature of the superconductivity. Such anisotropy has been reported in various two-dimensional superconducting systems, including Nd1-xEuxNiO2, infinite-layer lanthanum nickelates, LaMnO3/KTaO3, and NbSe2 [36, 37, 17, 38].
@@
Notably, the extrapolated in-plane upper critical field exceeds the Pauli paramagnetic limit, which is given by [39, 40]. Recent studies have demonstrated that strong spin-orbit coupling (SOC) in two-dimensional superconductors, can lead to out-of-plane spin locking and the formation of a uniaxial Ising spin texture[35, 41]. This spin-momentum locking suppresses Zeeman depairing, a key characteristic of Ising superconductivity, and can result in highly anisotropic superconducting behavior, with in-plane upper critical fields exceeding the Pauli limit. In this context, the enhanced in-plane upper critical field observed may be indicative of similar underlying physics.
@@
Figure 5A presents the current-voltage (I-V) characteristic curves as a function of temperature for Eu/p-BBO sample (Sample 2). From the I-V measurements, the critical currents () are determined and plotted in Figure 5B. The decreases with increasing temperature and eventually reaches zero at the transition temperature, as expected for a superconductor. We observe that the slope of I-V characteristics transitions from a normal ohmic state () to a non-linear power law () at the superconducting transition temperature. In 2D superconductors, this transition can be understood in the framework of the Berezinskii–Kosterlitz–Thouless (BKT) transition, where the emergence of nonlinear I-V characteristics in the superconducting state results from the current-driven unbinding of vortex-antivortex pairs, which are generated by thermal fluctuations at finite temperatures [42, 43]. The I-V characteristics are also plotted on a logarithmic scale in Figure 5C, highlighting the non-linearity more distinctly. To further analyze this transition, we extracted the exponent from the I-V curves and plot as a function of temperature in Figure 5D. The BKT transition temperature () is estimated to be around 5.34 K by interpolating to , which closely matches obtained from resistance measurements. Moreover, the temperature-dependent resistance near TBKT is expected to follow the Halperin-Nelson (HN) formula: R = R0 exp(-bt-1/2 ), where R0 and b are material-specific parameters and t = T/TBKT - 1 [44]. Figure S3 presents the fit based on the HN formula, yielding TBKT = 5.31 K, which agrees well with the value obtained from the I-V analysis. This agreement with BKT theory provides strong evidence for the two-dimensional (2D) nature of superconductivity in the heterostructures.
@@
Based on the structural transitions evidenced by the XRD and STEM measurements, the emergent superconductivity in the nominal Eu/p-BBO heterostructures arises from the formation of BaBi3 via the reduction of p-BBO. The structural results presented here rule out an electronic reconstruction at the interface between EuO and oxygen-deficient perovskite BaBiO3-x and/or electron doping of BaBiO3-x as the origin of superconductivity in nominal Eu/p-BBO heterostructures[20]. Studies of oxygen-deficient perovskite BaBiO3-x find an enhanced electronic band gap, suggesting that electron doping via oxygen vacancies may not stabilize superconductivity [45, 46]. We also consider the possibility of the reduction of p-BBO leading to the formation of an infinite-layered bismuthate structure analogous to the superconducting infinite-layer structures, which are also stabilized by the reduction of the perovskite phase[47, 48, 36, 49, 50]. First-principle calculations for an infinite-layer type BaBiO2 have recently been reported with a predicted Tc of 3.9–7.7 K under pressure and with spin-orbit coupling[51]. Here, superconductivity emerges on suppression of rotation of the oxygen sublattice at 10 GPa. However, the lattice parameters of the proposed structure are incompatible with our experimental results. It is, however, plausible that the layered infinite-layer phase forms as an intermediate phase and decomposes into BaO and BaBa3. For example, bulk stoichiometric NdNiO3 is known to be thermodynamically unstable and decomposes into Nd2O3, NiO, and Ni[52].

### v2

@@ Abstract
Oxygen-scavenging at oxide heterointerfaces has emerged as a powerful route for stabilizing metastable phases that exhibit interesting phenomena, including high-mobility two-dimensional electron gases and high Tc superconductivity. We investigate structural and chemical interactions at the heterointerface formed between Al or Eu and the charged-ordered insulator, BaBiO3, leading to emergent superconductivity at 6 K. A combination of X-ray diffraction and electron microscopy measurements shows that oxygen scavenging by the Eu and Al adlayers leads to the formation of superconducting intermetallic BaBi3 in nominal Eu/BaBiO3 and Al/BaBiO3 bilayers. Anisotropic magnetotransport measurements and current-voltage signatures of quasi two-dimensional superconductivity are observed. The mechanisms behind quasi-two-dimensional superconductivity and the role of disorder remain to be clarified. These findings highlight the potential for the use of in situ reduction of bismuthate heterostructures as a platform for stabilizing materials with exotic functional properties. Additionally, the strong spin-orbit coupling at the Bi sites may pave the way for the realization of high Tc topological superconductivity.
@@ I Introduction
Interest in Bi-rich superconductors continues due to the potential role of the strong spin-orbit coupling of Bi-6p electrons in enhancing superconductivity, and magnetic field-induced transitions from conventional s-wave to unconventional nodal d-wave superconductivity [9, 10]. Intermetallic ABi3 compounds (where A = Ba or Sr) have received attention because they exhibit superconductivity enhanced by spin-orbit coupling[11, 12, 9, 13]. Theoretical studies show that spin-orbit coupling suppresses strong Fermi nesting and increases electron-phonon coupling matrix elements associated with superconductivity[9]. At ambient pressure, BaBi3 crystallizes in a tetragonal structure (P4/mmm, a=5.06 , c=5.13 )[13, 12, 14, 15], featuring corner-sharing Bi6 octahedra (Figure 1B). Magnetotransport measurements on single crystals of BaBi3 and SrBi3 reveal superconducting transitions with onset temperatures of approximately 5.9 K and 5.6 K, respectively [12, 13]. Furthermore, scanning tunneling microscopy (STM) studies demonstrate the effect of symmetry breaking and disorder on the superconducting state in BaBi3 epitaxial films, grown in situ by molecular beam epitaxy[16]. However, to date, there have been no reports on the magnetotransport properties of thin BaBi3, due to its high air sensitivity and the availability of lattice matched substrates[13].
Here, we report the emergence of superconductivity at approximately 6 K in nominally Eu-capped p-BBO thin films grown via molecular beam epitaxy. X-ray diffraction measurements reveal that oxygen scavenging by the in-vacuo deposited metallic Eu or Al overlayer leads to the reduction of the monoclinic p-BBO layers and the decomposition of the reduced BBO films into tetragonal superconducting BaBi3 and rock-salt BaO. Temperature-dependent transport measurements, including angle-dependent magnetotransport and current-voltage characteristics, confirm the quasi-two-dimensional nature of the superconductivity. These findings highlight the potential of utilizing in situ reduction of bismuthate heterostructures as a platform for exploring novel quantum states. The strong spin-orbit coupling at Bi sites, may pave the way for the realization of high-temperature topological superconductivity, a long-sought phenomenon in perovskite-based systems[17, 18, 19].
@@
Synthesis
The samples were fabricated in a custom-oxide molecular beam epitaxy system. Bi, Ba, Eu, and Al were evaporated from effusion cells, with their growth rates calibrated by a quartz crystal monitor (QCM). p-BBO thin films of thicknesses ranging from 13 nm (30 unit cells) to 36 nm (90 unit cells) were grown on (001)-oriented SrTiO3 (STO) substrates at a substrate temperature of 585∘ C in an oxygen plasma with an oxygen partial pressure of 6.5x10-6 Torr and a plasma power of 250 W. Sharp reflection high-energy electron diffraction (RHEED) diffraction streaks observed during p-BBO growth shown in Figure 2A, are indicative of atomically flat p-BBO layers. The in-plane lattice parameter determined from the RHEED spacings is 4.35 , in agreement with bulk stoichiometric p-BBO [5]. Due to the large 11% lattice mismatch between p-BBO and STO, p-BBO grows relaxed via domain matching epitaxy.[20] For the nominal Eu/p-BBO samples, after growth, the p-BBO films were cooled to 500∘ C for the deposition of Eu metal in vacuo. RHEED patterns measured during Eu growth (Figure 2A) are consistent with the formation of cubic EuO from the topotactic reduction of the p-BBO layer [21, 22, 23]. The films were cooled in-vacuo and capped with 3 nm AlOx at room temperature to minimize oxidation of the EuO when exposed to ambient conditions[24]. A second set of samples was fabricated without Eu. Al was deposited at 500 ∘C in vacuum, after the depostion of p-BBO, and the samples were subsequently cooled to room temperature to form nominal Al/p-BBO bilayers.
@@
Figure 2B shows a comparison of specular scans along the (00L) direction for the nominal Eu/p-BBO sample and a single-component p-BBO film on STO. For the single-component p-BBO film, Bragg peaks are observed for the STO substrate and stoichiometric p-BBO. The calculated out-of-plane lattice constant for the p-BBO film is determined to be 4.36 in agreement with the measured pseudocubic values for the bulk monoclinic phase [5]. In contrast to the single-component p-BBO film, no p-BBO Bragg peaks were observed for the nominal Eu/p-BBO sample. Bragg peaks are observed at 32.316o and 34.997o corresponding to the (002) Bragg peaks of rock-salt BaO and face-centered tetragonal BaBi3, respectively. The calculated out-of-plane lattice constants are 5.52 and 5.11 in good agreement with the bulk lattice constants for BaO (5.53 ) and BaBi3 (5.13 ) [25, 12]. The lattice constant of NaCl-type EuO is 5.1439 ,[26] hence, the EuO (002) Bragg peak will overlap with the BaBi3 (002) peak.
To confirm the formation of BaBi3, XRD measurements were carried out on samples with Al metal deposited directly on a p-BBO film at 500∘C. We find that the Al metal effectively reduces the p-BBO film to form BaBi3 and BaO as evidenced by XRD measurements in Figure 2B. Reciprocal space maps (RSMs) around the STO (102), BaO (113), and BaBi3 (113) Bragg peaks are shown in Figure S1 in the Supplementary Materials[27]. The BaBi3 is oriented with the BaBi3 [0 0 1] axis parallel to the SrTiO3 [0 0 1] axis and the BaBi3 [1 1 0]//SrTiO3 [1 0 0] axis. The lattice constants are determined to be a= 5.21 and c=5.15 in agreement with bulk values of -BaBi3 [14]. The RSM confirms that the BaBi3 is fully relaxed while the BaO is lattice-matched to the STO substrate. The lattice relaxation of BaBi3 is expected due to the 8% mismatch between the (110) planes with and STO (.
Next, we investigate the effect of the ratio of the thickness of the Al and BaBiO3 on the formation of the BaBi3 phase. We compare the X-ray diffraction patterns of 45 monolayers (ML) (1 ML p-BBO = 4.35 ) of BaBiO3 capped with 0, 10, and 45 ML of Al in Figure S2 in the Supplementary Materials [27]. For the 10 ML Al/45 ML BaBiO3 sample, the perovskite BaBiO3 shifts to lower angles consistent with the formation of partially reduced BaBiO3-x and an increase of the c-axis lattice parameter [28]. The perovskite phase is suppressed in the 10 ML Al/45 ML BaBiO3 sample, confirming that further reduction of the perovskite BaBiO3 phase leads to decomposition in BaO and BaBi3.
@@
Scanning transmission electron microscopy (STEM) measurements were performed on the films to further investigate the structural properties of the heterostructures. Figure 2C shows a high-angle annular dark field (HAADF) STEM image of a cross-section of a nominal Eu/p-BBO film on STO. Sharp interfaces are observed between the nominal layers, indicative of insignificant intermixing. The energy-dispersive X-ray spectroscopy (EDX) profiles of the section are shown in Figure 2D. The EDX maps reveal significant chemical segregation in the nominal p-BBO region. The nominal Eu and STO regions are oxygen-rich, while the nominal p-BBO section is oxygen-poor, consistent with the reduction of p-BBO by the Eu adlayer. Ba-rich regions are observed at the nominal p-BBO/STO and Eu/p-BBO interfaces. In the nominal p-BBO region, we observe lateral phase-separated Ba and Bi-rich regions, as shown in Figure S3 in the Supplementary Materials[27]. The oxygen content in the Ba-rich regions is significantly higher than in the Bi-rich sections, consistent with the BaO and BaBi3 Bragg peaks observed in the XRD measurements.
Transport Measurements
Next, we investigate the transport properties of nominal Eu/p-BBO heterostructures from 1.8 - 300 K. Figure 3A shows the temperature-dependent resistance of a nominal Eu (16 nm)/p-BBO (36 nm) heterostructure (Sample 1). As the temperature decreases to 7 K, the resistance decreases, exhibiting characteristic metallic behavior. The negative curvature of the resistivity between 50-80 K is consistent with the transport properties of bulk BaBi3 and may be associated with two-band conductivity [12, 29, 30]. Between 25 K and 6 K, the resistance exhibits a linear temperature dependence. This linearity is more clearly demonstrated by the linear fitting of the resistance vs. temperature data shown in Figure S4 (A). This behavior is indicative of strange-metal characteristics commonly observed in unconventional superconductors [31].
Upon further cooling, the sample undergoes a transition to a superconducting state with a narrow transition width of approximately 0.133 K, reaching a zero resistance state at 5.77 K (). The onset superconducting temperature (defined as the temperature at which the resistance drops to 90% of its normal state value), is 6.03 K, significantly higher than the previously reported value of 3.5 K for EuO/BaBiO3 heterostructures[21] but in agreement with the reported values for bulk BaBi3[12, 13, 9, 14, 15]. Above Tc, for B = 4T, linearity is suppressed along with superconductivity, consistent with previous reports [12, 32, 33]. We performed a T2 fit, characteristic of Fermi-liquid behavior, to the resistivity curve for B = 4 T as shown in Figure S4 (B). A clear deviation from T2 behavior is observed above 10K, in agreement with transport measurements for bulk BaBi3 [12]. These results combined with XRD results suggest the emergence of superconductivity is a result of the formation of BaBi3.
The inset of Figure 3A shows the temperature-dependent resistance of Sample 1 with a magnetic field (), applied out of the plane. is suppressed as increases, further confirming superconductivity. Magnetoresistance curves with at different temperatures close to the superconducting transition are plotted in Figure 3B. In the normal state above Tc, the resistance exhibits a quadratic dependence on the applied magnetic field, similar to conventional metals[34]. Below Tc, a resistance peak appears above the critical magnetic field, suggesting disorder arising from granularity[35]. As the BaBi3 phase emerges from the decomposition of p-BBO into BaO and BaBi3, inhomogeneity and defect phases are likely to form giving rise to inhomogeneous superconductivity.
@@
The temperature-dependent Hall resistance exhibits a linear dependence on the magnetic field, B, as shown in Inset Figure 3C. A change in slope and, consequently, the sign of the Hall coefficient occurs at 65 K. This change indicates a shift from electron-dominated conduction at low temperatures to hole-dominated conduction at higher temperatures, suggesting the coexistence of both electron and hole pockets at the Fermi level, in agreement with previously reported theoretical predictions [9]. The sheet carrier density and mobility at 10 K are 1.6 1016 cm-2 and 1.2 cm2/Vs, respectively. These carrier density values are consistent with previous observations[21]. The carrier concentrations are 3-4 orders of magnitude higher than observed for the two-dimensional electron gas at SrTiO3 [36, 37] and KTaO3 [38, 39, 40, 41] interfaces. Furthermore, Figure 3D compares the resistance as a function of the nominal p-BBO thickness. The number of Eu monolayers is tuned to match the p-BBO thicknesses. The transition temperatures are comparable and independent of the nominal p-BBO thicknesses.
@@
The critical fields for both out-of-plane, , and in-plane configurations were determined using the 90% and 50% of normal resistance criterion. and are plotted as a function of temperature in Figure 4D. To determine the upper critical fields at T = 0 K, the data were fitted using the two-dimensional Ginzburg-Landau (GL) model for the out-of-plane configuration and in-plane configuration [42]:
and
@@
where and represent the extrapolated upper critical fields at 0 K. The extracted values of using the 90% and 50% criteria are 5.54 T and 4.66 T, respectively, both exceeding the reported value of 2.2 T for bulk BaBi3 single crystals [12, 43]. As shown in Figure 4D, the temperature dependence of the in-plane upper critical field determined using the 90% criterion deviates from the two-dimensional (2D) Ginzburg-Landau (GL) model, whereas determined by the 50% criterion follows the expected 2D-GL behavior. The extrapolated value of (50%) is 7.49 T. These critical fields satisfy the GL relations and , where is the magnetic-flux quantum, is the coherence length, and is the effective superconducting thickness. Using the extracted upper critical-field values (50% criterion), we obtain nm, smaller than nm, consistent with quasi-two-dimensional behavior reported for Eu/BBO heterostructures [21]. The difference between the in-plane critical fields derived from the 50% and 90% criteria originates from a multistage superconducting transition feature appearing near 5 T (Figure 4C). The anisotropy of in-plane and out-of-plane upper critical fields has been reported in various two-dimensional superconducting systems, including Nd1-xEuxNiO2, infinite-layer lanthanum nickelates, LaMnO3/KTaO3, and NbSe2 [44, 45, 17, 46].
@@
Recent studies have demonstrated that strong spin-orbit coupling (SOC) in two-dimensional superconductors can lead to out-of-plane spin locking and the formation of a uniaxial Ising spin texture[43, 32]. This spin-momentum locking suppresses Zeeman depairing, a key characteristic of Ising superconductivity, and can result in highly anisotropic superconducting behavior, with in-plane upper critical fields exceeding the Pauli limit. In contrast, Rashba-type SOC leads to in-plane spin locking, which enhances in-plane to approximately BP[47]. Furthermore, spin-orbit scattering can randomize electron spins, reducing paramagnetic limiting effects and thereby increasing [48]. Notably, the Pauli paramagnetic limit, given by , is about 10.6 T for the BaBi3 films [49, 50]. While the extracted value of (50 %) = 7.49 T is below this Pauli limit, Figure 4C shows that the superconducting state is not fully suppressed for 8 T and may approach the Pauli limit at 0 K.
@@
Figure 5A presents the current-voltage (I-V) characteristic curves as a function of temperature for Sample 2. From the I-V measurements, the critical currents () are determined and plotted in Figure 5B. decreases with increasing temperature and eventually reaches zero at the transition temperature, as expected for a superconductor. We observe that the slope of I-V characteristics transitions from a normal ohmic state () to a non-linear power law () at the superconducting transition temperature. In 2D superconductors, this transition can be understood in the framework of the Berezinskii-Kosterlitz-Thouless (BKT) transition, where the emergence of nonlinear I-V characteristics in the superconducting state results from the current-driven unbinding of vortex-antivortex pairs, which are generated by thermal fluctuations at finite temperatures [51, 52]. The I-V characteristics are also plotted on a logarithmic scale in Figure 5C, highlighting the non-linearity more distinctly. To further analyze this transition, we extracted the exponent from the I-V curves and plot as a function of temperature in Figure 5D. The BKT transition temperature () is estimated to be around 5.34 K by interpolating to , which closely matches obtained from resistance measurements. Moreover, the temperature-dependent resistance near TBKT is expected to follow the Halperin-Nelson (HN) formula: R = R0 exp(-bt-1/2 ), where R0 and b are material-specific parameters and t = T/TBKT - 1 [53]. Figure S5 (Supplementary Materials[27]) presents the fit based on the HN formula, yielding TBKT = 5.31 K, which agrees well with the value obtained from the I-V analysis. The BKT transition is also measured in a third sample in Figure S6 in the Supplementary Materials.[27] This agreement with BKT theory is consistent with two-dimensional (2D) nature of superconductivity in the heterostructures.
To examine the role of the capping layer, we performed magnetotransport measurements on nominal Al/p-BBO heterostructures. Figure 6A presents the temperature-dependent resistance, which shows superconducting behaviour with a transition temperature of 5.75 K, comparable to that observed in Eu/p-BBO. Magnetoresistance curves measured with the out-of-plane perpendicular magnetic field (B⟂) at various temperatures are shown in Figure 6B. Figure 6C shows the resistance as a function of temperature for different values of with a decrease in observed with increasing . The out-of-plane critical field, , extracted using the 50% normal resistance criterion, is plotted as a function of temperature in Figure 6D. The value of extracted by fitting the 2D-GL equation for the out-of-plane configuration is 5.52 T, which is close to the values observed in the Eu/p-BBO samples. These results indicate that the choice of capping layer has little influence on the superconducting properties of BaBi3.
@@
Based on the structural transitions evidenced by the XRD and STEM measurements, the emergent superconductivity in the nominal Eu/p-BBO and Al/p-BBO heterostructures arises from the formation of BaBi3 via the reduction of p-BBO. The structural results presented here rule out an electronic reconstruction at the interface between EuO and oxygen-deficient perovskite BaBiO3-x and/or electron doping of BaBiO3-x as the origin of superconductivity in nominal Eu/p-BBO heterostructures[21]. Studies of oxygen-deficient perovskite BaBiO3-x find an enhanced electronic band gap, suggesting that electron doping via oxygen vacancies may not stabilize superconductivity [28, 54]. We also consider the possibility of the reduction of p-BBO leading to the formation of an infinite-layered bismuthate structure analogous to the superconducting infinite-layer structures, which are also stabilized by the reduction of the perovskite phase[55, 56, 44, 57, 58]. First-principle calculations for an infinite-layer type BaBiO2 have recently been reported with a predicted Tc of 3.9 - 7.7 K under pressure and with spin-orbit coupling[59]. Here, superconductivity emerges on suppression of rotation of the oxygen sublattice at 10 GPa. However, the lattice parameters of the proposed structure are incompatible with our experimental results. It is however, plausible that the layered infinite-layer phase forms as an intermediate phase and decomposes into BaO and BaBi3. For example, bulk stoichiometric NdNiO2 is known to be thermodynamically unstable and decomposes into Nd2O3, NiO, and Ni[60].
The anisotropic critical magnetic fields and the presence of a BKT transition are consistent with quasi-two-dimensional superconductivity.[21] However, the crystal structure of BaBi3 is inherently three-dimensional, and the nominal thicknesses of the p-BBO films (13-36 nm) exceed the superconducting coherence length, nm. In addition, structural and chemical disorder arising from phase separation between BaO and BaBi3 may lead to spatially inhomogeneous superconductivity. Another relevant length scale for assessing dimensionality is the superconducting London penetration depth, . For bulk BaBi3, nm, which is an order of magnitude larger than the film thickness .[15, 12] Therefore, in the limit , the superconductivity can be effectively regarded as two-dimensional. Future studies should aim to unravel the microscopic mechanisms governing the quasi-two-dimensional superconducting state and quantify the extent to which disorder modulates these phenomena.